You have an order for phenytoin elixir 100 mg PO. The label on the bottle reads 125 mg per 5mL. How many ml will you give?
Answer #1

Ordered dose= 100 mg

5 millilitres= 125 milligrams

So 1 ml = 125/5 = 5 milligrams

1 ml contains 5 milligrams

Thus , 100 milligram =

ratio method

5 mg/ 1 ml= 100 mg/x ml.( 5 mg in 1 ml = 100 mg in x ml)

5x= 100

X= 100/5= 50

100 mg = 50 ml

prescribed dose in ml= 50 ml.(100 mg)